NetCall - Release 2

Poll-handler for Fido-Net-Systems Version 2.0a

(Last change: 14.02.95)

by Alexander Stirmlinger

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994-1995 Alexander Stirmlinger

NetCall is shareware. All rights of the program, files and documents belong completely to the author. The archive and its files (program, documents and the keyfile) must not be changed in any way. No files may be added or deleted to/from the archive.

You may charge a fee to recover distribution costs. The fee for diskette distribution may not be more than the cost to obtain a public domain diskette from Fred Fish.

1 General informations

1.1 Features of NetCall

• Simplifies the treatment of file-requests and crash-mails (pending jobs are displayed in a listview-gadget).

• Enables automatic polling on all lines of the destination system (if one line is BUSY, NetCall will automatically dial up the next line).

• Searches (even when your mailer polls already) for other lines of the destination system in the nodelist (crashing or frequesting).

• Supports/simplifies multi-bossing.

• Simplifies polling on a line of your boss, which is not listed in the nodelist.

• Hides your own password when crashing or frequesting.

• Supports automatical operation (without user-interaction).

• Font-sensitive GUI.

• Opens public-screen if desired.

• User-defineable menus (user-commands).

• Extensive ways to configure.

• Many shell-arguments.

• Localized (OS2.1 and better).

• Iconify-function (including QUIET-mode).

1.2 Copyright

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994-1995 Alexander Stirmlinger

NetCall is shareware. All rights of the program, files and documents belong completely to the author. The archive and its files (program, documents and the keyfile) must not be changed in any way. No files may be added or deleted to/from the archive.

The keyfile is personalized to its owner and must not be changed, saled, hired or lended.

1.3 Warranty

NetCall is provided "AS-IS". I can’t guarantee the expected work of the program.

Though, I’m always anxious to remove all bugs and misbehaviors.

Reports, hints and suggestions are always welcome. (See also Chapter 4 [Author info], Page 11.)

  THERE IS NO WARRANTY FOR THE PROGRAMS, TO THE EXTENT PERMITTED
BY APPLICABLE LAW.  EXCEPT WHEN OTHERWISE STATED IN WRITING THE
COPYRIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAMS "AS
IS" WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED,
INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF
MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE.  THE
ENTIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE PROGRAMS IS
WITH YOU.  SHOULD THE PROGRAMS PROVE DEFECTIVE, YOU ASSUME THE
COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.
  IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN
WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY
REDISTRIBUTE THE PROGRAMS AS PERMITTED ABOVE, BE LIABLE TO YOU
FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR
CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE
THE PROGRAMS (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA
BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD
PARTIES OR A FAILURE OF THE PROGRAMS TO OPERATE WITH ANY OTHER
PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED
OF THE POSSIBILITY OF SUCH DAMAGES.

1.4 Restrictions of the unregistered version

• Nerving requester

• Nerving wait-function (appears randomly)

2 Installation

2.1 Assumptions to the hardware

NetCall does not make any assumptions to the hardware. NetCall runs on every Amiga with at least 512 KB RAM.

2.2 Assumptions to the software

NetCall runs on Amigas equipped with AmigaOS 2.04 or later. AmigaOS 3.x is, of course, supported, too.

Additionally, you need TrapDoor^{1} (not included in the NetCall-archive) as a FidoNet-mailer. Other mailers may only be used, if they use the same shell-arguments and a compatible logfile (NetCall needs to scan the TrapDoor-logfile).

2.3 Installer

The installation is easily done by Commodore-Amiga’s Installer. Just double-click on one of the install-icons (english or german language). All needed files will be copied to your hard-disk.

After having installed NetCall, you’ll have to configure NetCall to your wishes and to your FidoNet-software. Please read the configuration-chapter of this document for further information. See also Chapter 3 [Configuration], Page 7.

Installer and Installer project icon (c) Copyright 1991-93 Commodore-Amiga, Inc. All Rights Reserved. Reproduced and distributed under license from Commodore.

INSTALLER SOFTWARE IS PROVIDED "AS-IS" AND SUBJECT TO CHANGE; NO WAR-RANTIES ARE MADE. ALL USE IS AT YOUR OWN RISK. NO LIABILITY OR RESPONSI-BILITY IS ASSUMED.

1 TrapDoor is copyright \begin{matrix}c \\ ◯ \end{matrix} Maximilian Hantsch and Martin Laubach

2.4 Execution from a shell

The following arguments may be passed if NetCall is executed from a shell:

POLLID,CALL/K,CONFIG/K,OUTBOUND/K,PUBSCREENNAME/K,NOREQ/S,
NOPOLL/S,NOCRASH/S,NOFREQ/S,ALLACTIVE/S,OPENBIG/S,TERMREQS/S

POLLID PollID(s)

This is the ID from your bosses, where you want NetCall to poll. The IDs given here must correspond with the BOSS-entries in your ‘NetCall.cfg’ (case-sensitive). You may pass a valid AmigaDOS-pattern as an argument here (‘POLLID (BBS1|BBS2)’ would cause NetCall to poll at node BBS1 and BBS2).

CALL Node/Phone Node- or phone-number where you want NetCall to poll

(example: ‘CALL 2:246/1545’).

CONFIG Config-file Use the given configuration-file rather than the normal configuration.

(example: ‘CONFIG Work:NetCall/Config2.cfg’).

OUTBOUND Outbound-path Scan this OUTBOUND-directory rather than the directory specified in the configurationfile.

PUBSCREENNAME public-screen

Force to open NetCall’s and the mailer’s windows on this public-screen.

NOREQ

Use stdIO for error-output rather than a system-requester.

NOPOLL

Do not pick up entries where NetCall would just poll (in other words: exclude your bosses).

NOCRASH

Exclude entries where NetCall would crash mails to.

NOFREQ

Exclude entries where NetCall would file-request from.

ALLACTIVE

Mark all entries as active immediately after NetCall gets started. NetCall will start to poll immediately.

OPENBIG

Always open NetCall’s big window (two list-view-gadgets, one text-gadget).

TERMREQS

Forces NetCall to terminate requesters (even error-requesters!) automatically. This option allows automatical operation without user-interaction.

2.5 Execution from workbench

NetCall can easily be started by double-clicking on it’s icon.

The following arguments may be passed as tool-types in the program icon:

ARGS = <shell-args> All arguments, which can be passed as shell-arguments, can be passed in the ARGStool-type, too. See also Section 2.4 [Execution from a shell], Page 4.

TOOLPRI = <priority (-127 to 128)> Taskpriority of NetCall.

6 NetCall 2.0a

3 Configuration

3.1 The configuration-file NetCall.cfg

NetCall will look for its configuration-file in the following directories:

1. Mail: (if assigned)

2. The current directory

3. The directory, where the environment-variable NETCALLCONFIG points to

4. Shell-argument CONFIG (See also Section 2.4 [Execution from a shell], Page 4.)

The configuration-file of NetCall may/must include the following parameters:

OUTBOUND "Path"

OUTBOUND-directory of your mailer (example: ‘Mail:Outbound’).

NODELIST "Path"

Directory where your nodelist can be found (example: ‘Mail:Nodelist’). If you don’t have a nodelist, delete this line from your configuration-file or mark it out as a comment.

MAILER "Path/Mailer"

TrapDoor-mailer (example: ‘Mail:Bin/TrapDoor’).

ARGAPPEND "Mailer shell-arguments" Arguments specified here, will be passed to the mailer, directly (example: ‘BACKGROUND’).

ARGAPPENDICONIFIED "Mailer shell-Arguments" Additional shell-arguments passed to the mailer, when NetCall operates in the iconified mode (example: ‘QUIET’ would cause TrapDoor to open no windows, when NetCall is iconified).

TDLOGFILE "Path/TrapDoor-LogFile"

LogFile of TrapDoor (example: ‘Mail:Logs/TrapDoor.log’).

NETCALLSTART "Shell-command"

This shell-command gets executed at NetCall’s startup.

NETCALLEND "Shell-command"

This shell-command gets executed at NetCall’s shutdown.

HELPCOMMAND "Shell-command"

This shell-command gets executed if you press the <HELP>-key when the NetCallwindow is active.

(‘SYS:Utilities/AmigaGuide HELP:english/NetCall.guide PUBSCREEN NETCALL’).
REDIALDELAY seconds

Time that should delay if all lines are BUSY.

MAXRETRIES retries

After having reached this amount of retries, NetCall will automatically deactivate all active list-entries. (Use ‘MAXRETRIES 0’, if you don’t want automatical deactivation.

MULTILINE

This keyword will cause NetCall to search for other nodes of your destination system(s) in the nodelist. The search will be performed by a second task, so no delay of the polling-task should be noticeable. Use NOMULTILINE if you don’t want this option.

AUTOQUIT

NetCall will be automatically terminated as soon as all jobs are processed. Use NO-AUTOQUIT if you don’t want this option.

MAILERABORT (ABORTNETCALL|CONTINUE|REQUEST)

If you quit your mailer (user-break, <ESC>), NetCall may:

• be quitted, too (MAILERABORT ABORTNETCALL)

• poll at the next node (MAILERABORT CONTINUE)

• open a requester (MAILERABORT REQUEST)

OPENPUBSCREEN "Screenname" Width Height Depth ScreenModeID Force NetCall to open its own public-screen. All values must be given in decimal format. To specifiy the ScreenModeID, you may use the ScreenModeViewer-Utility

(example: ‘NETCALL 640 256 2 167936’).
PUBSCREEN "Screenname"

Open NetCall on this public-screen. If NetCall can’t find this screen to open on, NetCall will open its window on your workbench. If you wish NetCall to open its own public-screen (OPENPUBSCREEN) you must specifiy the same name at OPEN-PUBSCREEN and PUBSCREEN (example: ‘NETCALL’):

PUBSCREENMAILER "Screenname"

You can optionally force the mailer to open its windows on this public-screen (example:

‘Workbench’).

WINDOW X/Y/Width/Height Window-position/-size (big NetCall-window, incl. list-view-gadgets and text-gadget)

(example: ‘30/30/400/180’).

WINDOWZOOM X/Y/Width/Height Window-position/-size of the zoomed (little) NetCall-window. If you specifiy AUTO rather than a decimal value as the height, NetCall will automatically adapt its windowheight to your actual font-size (example: ‘30/30/400/AUTO’).

WINDOWACTIVATE

Activate the NetCall-window when opened (use NOWINDOWACTIVATE if you don’t want to activate).

WINDOWFONTNAME "Fontname"

The Font that should be used for the list-view-gadgets (example: ‘topaz.font’).

WINDOWFONTSIZE Height

Font-size used for the list-view-gadgets (example: ‘8’).

LVITEMHEIGHT Pixel

Space between the list-view-entries (given in pixel) (example: ‘1’).

OPENBIG

If you specify OPENBIG, NetCall will always open its big window, no care if there are file-requests or crash-mails in your OUTBOUND-directory. Use NOOPENBIG if you don’t want this option.

SEARCHOPTS 5/6

Search-options (for MULTILINE). Leave it be...

SEARCHPATTERN "Pattern"

Pick up all nodes with this fido-flags to the ToDo-list. A valid AmigaDOS pattern can be given here. If you specify ‘~(#?ISDN#?)’ NetCall will pick up all fido-nodes without ISDN. If you don’t specify MULTILINE, this keyword will have no effect.

COMMAND "Name of the menu-entry" "Shell-command" ASYNC With this keyword you are abled to define your own menu-entries in the NetCallmenustrip. You may specify any number of COMMAND-keywords here. If you add ASYNC, NetCall will launch the command as an seperate process. (example: ‘Test1

SYS:Prefs/Time’).

BOSS ID(*) Node/Phone Point Password Retries With the BOSS-keyword you can define all lines of your own fido-bosses. You may specify any number of BOSS-keywords. By using several different IDs you can configure your system for multi-bossing. Entries with a * behind the ID will be displayed in the NetCall-window (list-view-gadget). You may use different amounts of retries for each line of your boss. You can specify phone-numbers rather than node-numbers if your boss has unlisted lines. Example:

      BOSS  SO*   2:246/2248   2:246/2248.20  test  3
      BOSS  SO    2:246/2249   2:246/2248.20  test
      BOSS  SWP*  2:246/1545   2:246/1545.11  test
      BOSS  SWP   2:246/1546   2:246/1545.11  test
      BOSS  SWP   2:246/1547   2:246/1545.11  test  3
      BOSS  SWP   07131167890  2:246/1545.11  test

3.2 The configuration of TrapDoor

You should set the amount of RETRIES in your ‘TrapDoor.cfg’ to 0 (‘RETRIES 0’), because NetCall will do this from now on.

3.3 The usage of NetCall with Spot

To use NetCall with Spot^{1} as optimized as possible, you should configure Spot as follows: Execute NetCall only from your first boss-entry in the Spot-configuration (Settings -> Sy-

stem...). Enable the asynchronous-button.

   NetCall >NIL: OPENBIG

All other NetCall-entries of Spot should be cleared. If you press the NetCall-button of Spot, NetCall will be started. If you have multiple bosses, Spot will no longer ask you where to poll. You can choose this

directly in the NetCall-window.

3.4 The usage of NetCall with UMS

If you want to use NetCall with UMS^{2} you don’t have to change anything in your ‘ums.config’. NetCall can easily be started by double-clicking its icon on the workbench or executing it in a

shell.

1 Spot is copyright \begin{matrix}c \\ ◯ \end{matrix} Nico François 2 UMS is copyright \begin{matrix}c \\ ◯ \end{matrix} Martin Horneffer

4 Author information

You can always reach me as follows:

     Alexander Stirmlinger      Home: Weinbergweg 10
                                      D-74831 Gundelsheim
                           Phone/Fax: ++49 - 62 69 - 89 92
                             FidoNet: 2:246/1545.11

12 NetCall 2.0a

5 Acknowledgements

I’d wish to thank the following people for their bug-reports, their many suggestions and their endless support during the beta-test:

Martin Biallas
Frank Dürring
Klaus Gessner
 Stefan Hirth
Manfred Hoffmann
Peter Jachmann
Gerald Kleesiek
 Markus Mayer
  Sven Meyer
Werner Senden

14 NetCall 2.0a

Appendix A How to register

If you want to use NetCall regularly, you have to register! If you register, you can use NetCall without any shareware-reminders. See also Section 1.4 [Restrictions of the unregistered version], Page 2.

Once registered, you are able to use all forthcoming versions of NetCall without sharewarereminders; your keyfile will be valid forever.

NetCall is shareware and costs 10,- DM (10.- US$). The following data is required for registration: your name, street, ZIP-code, city, telephone-

number and your email-address(es).

You may pay the share as follows:

1. Euro-cheque: DM (Deutsche Mark), payable to Alexander Stirmlinger.

2. Wire bank transfer: Please give NetCall as a comment. Please send me your address etc. via email or mail.

3. Cash: DM or US$ - at your own risk.

If you wish to get your keyfile PGP-encoded, send me your PGP-public first. If you want to get your keyfile by mail, please send me a blank disk and a self-addressed and

stamped envelope.

Bank: Kreissparkasse Heilbronn Account : 190 3780 BLZ : 620 500 00

Thank you for supporting shareware!

16 NetCall 2.0a

Appendix B Support/Updates

You can always file-request or download the most recent version of NetCall at the CyberDyne System, Heilbronn, Germany (magic: NETCALL).

Registered users will receive release-notes as carbon-copies via email, if new NetCall-releases are available.

CyberDyne System (24h online):
Node 1: +49-7131-176169 * 2:246/1545 * 1200 - 16800 BPS
        V.32B V.42B ZYX
Node 2: +49-7131-161614 * 2:246/1546 * 1200 - 14400 BPS
        V.32B V.42B HST
Node 3: +49-7131-167890 * 2:246/1547 * 1200 - 28800 BPS
        V.32B V.42B V.FC
Node 4: +49-7131-950371 * 2:246/1548 * 64000 BPS BLASTER
        X.75 V.110 ISDN
Node 5: +49-7131-950372 * 2:246/1549 * 64000 BPS ELINK
        X.75 V.110 ISDN

18 NetCall 2.0a

Appendix C History

C.1 NetCall - Version 2.0

• First public release.

C.2 NetCall - Version 2.0a

• English documents included.

• New config-variable: MAXRETRIES.

• Minor ALLACTIVE-bug fixed.

• New shell-argument: TERMREQS. Supports automatical operation.

20 NetCall 2.0a Index 21

Index

A

Assumptions to the hardware . . . . . . . . . . . . . . . . . . . . . . . 3 Assumptions to the software . . . . . . . . . . . . . . . . . . . . . . . . 3 Author information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

C

Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Copyright. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

E

Execution from workbench . . . . . . . . . . . . . . . . . . . . . . . . . . 5

F

Features of NetCall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 FidoNet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

G

General informations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

H

History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

I

Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Installer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

K

Keyfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

N

NetCall.cfg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 No warranty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

P

Paramenter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 5 PGP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

R

Restrictions of the unreg. version . . . . . . . . . . . . . . . . . . . 2

S

ShareWare. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Shell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Shell-arguments. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Spot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

T

Tool-types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 TrapDoor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 10 TrapDoor.cfg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

U

UMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Updates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15, 17

W

Warranty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Workbench . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

22 NetCall 2.0a

Table of Contents

1 General informations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Features of NetCall. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 Warranty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.4 Restrictions of the unregistered version . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2.1 Assumptions to the hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Assumptions to the software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.3 Installer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.4 Execution from a shell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.5 Execution from workbench . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3 Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

3.1 The configuration-file NetCall.cfg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.2 The configuration of TrapDoor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.3 The usage of NetCall with Spot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.4 The usage of NetCall with UMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

4 Author information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5 Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Appendix A How to register . . . . . . . . . . . . . . . . . . . . . . . . . 15 Appendix B Support/Updates . . . . . . . . . . . . . . . . . . . . . . . 17 Appendix C History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

C.1 NetCall - Version 2.0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 C.2 NetCall - Version 2.0a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

ii NetCall 2.0a